# FactSelfCheck: Fact-Level Black-Box Hallucination Detection for LLMs

### v1

@@ Abstract
Large Language Models (LLMs) frequently generate hallucinated content, posing significant challenges for applications where factuality is crucial. While existing hallucination detection methods typically operate at the sentence level or passage level, we propose FactSelfCheck, a novel black-box sampling-based method that enables fine-grained fact-level detection. Our approach represents text as knowledge graphs consisting of facts in the form of triples. Through analyzing factual consistency across multiple LLM responses, we compute fine-grained hallucination scores without requiring external resources or training data. Our evaluation demonstrates that FactSelfCheck performs competitively with leading sampling-based methods while providing more detailed insights. Most notably, our fact-level approach significantly improves hallucination correction, achieving a $35\%$ increase in factual content compared to the baseline, while sentence-level SelfCheckGPT yields only an 8% improvement. The granular nature of our detection enables more precise identification and correction of hallucinated content.
@@ 1 Introduction
Figure 1: The pipeline of FactSelfCheck in two variants. For response $p$, entities $\mathcal{E}_{p}$ and relations $\mathcal{R}_{p}$ are extracted, followed by the construction of knowledge graphs $KG_{p}$, for which hallucination scores $\mathcal{H}_{\text{fact}}$ are calculated. Samples’ entities $\mathcal{E}_{S}$ and relations $\mathcal{R}_{S}$ are created by merging $\mathcal{E}_{p}$ and $\mathcal{R}_{p}$ with entities and relations from $KG_{p}$. For each sample $s$, the knowledge graph $KG_{s}$ is extracted. FactSelfCheck-KG assesses the consistency between a fact and all $KG_{s}$. FactSelfCheck-Text assesses the consistency between a fact and all $s$ directly. To obtain sentence-level $\mathcal{H}_{\text{sentence}}$ and passage-level $\mathcal{H}_{\text{passage}}$ scores, fact-level scores are aggregated, as indicated by dashed arrows.
@@
Our granular approach is motivated by two key observations. First, a single sentence or text passage can contain multiple facts, with the number of facts varying significantly across sentences, contexts, and domains. This variability makes it challenging to identify hallucinated aspects of generated output precisely when using sentence-level or passage-level detection. Second, false information can be dispersed throughout a text, as a single fact may appear across multiple sentences. That can mislead the sentence-level detection, as the factuality of a sentence is dependent on the previous sentences Zhang et al. (2024). These challenges become particularly acute when analyzing long texts generated by LLMs. Fine-grained fact-level detection provides a more precise understanding of text factuality than a sentence or passage analysis. It enables better assessment of content reliability and, as we show later, more effective factuality correction.
@@
We propose FactSelfCheck, a black-box method for fact-level hallucination detection, meaning it does not require access to the model’s internal parameters. This design choice makes our approach universally applicable across any LLM, including closed models, like GPT OpenAI et al. (2024). Following a sampling-based detection paradigm, introduced by Manakul et al., 2023, our method utilizes multiple response generations and analyzes the factual consistency of extracted facts across these samples. This paradigm is based on the phenomenon that factual information remains largely consistent across different generations, while hallucinated content tends to vary or contradict itself between samples Manakul et al. (2023); Wang et al. (2023). This way, we can effectively identify hallucinated facts without relying on external resources (zero-resource) or access to the model’s internal parameters (black-box). Moreover, our method is non-parametric, as it does not require any training, making it easy to implement in any domain without the need for training data. The FactSelfCheck pipeline consists of three main steps: knowledge graph extraction, which extracts sets of facts from the initial response and samples; fact-level hallucination scoring; and calculating sentence-level and passage-level scores by aggregating fact-level scores.
@@
We evaluated our method using the WikiBio GPT-3 Hallucination Dataset Manakul et al. (2023), aggregating fact-level scores into sentence-level and passage-level scores. Our approach achieves performance comparable to leading sampling-based methods while providing more detailed information about hallucinations. Additionally, we demonstrate that our fact-level approach significantly improves hallucination correction. Compared to a baseline, providing incorrect facts to the correction method leads to a $35\%$ increase in factual content, while passing incorrect sentences leads to an $8\%$ increase.
@@
    The novel black-box sampling-based method for fact-level hallucination detection – FactSelfCheck. It enables fine-grained hallucination detection in LLM responses without requiring training data or external resources, as it is both non-parametric and zero-resource.
@@
    The effective approaches for measuring factual consistency across multiple samples: FactSelfCheck-KG using knowledge graph comparisons and FactSelfCheck-Text using direct text comparison.
@@
Our code is available on GitHub  [^1] https://github.com/graphml-lab-pwr/FactSelfCheck license: CC BY-SA 4.0.
@@ 2 Related work
Black-box approaches operate without access to the model’s internal states and aim to detect hallucinations based solely on the text generated by LLMs. Some of these methods use external resources to collect evidence (Min et al., 2023; Chern et al., 2023). Others leverage LLMs to detect hallucinations like CoVe Dhuliawala et al. (2024), which utilizes the chain-of-thought paradigm for detection. Another category is sampling-based methods, such as SelfCheckGPT (Manakul et al., 2023), which evaluate factuality by generating multiple responses (stochastic samples) and assessing consistency between the original response and these samples.
@@ 3 Method
We propose FactSelfCheck, a black-box sampling-based method for fact-level hallucination detection, as illustrated in Figure 1. We list used prompts in Appendix F.
@@ 3.1 Notation
Our objective is to compute a fact-level hallucination score $\mathcal{H}_{\text{fact}}$ for each fact $f$ in $KG_{p}$. Subsequently, to facilitate comparisons with other methods, we aggregate these scores to obtain a sentence-level hallucination score $\mathcal{H}_{\text{sentence}}$ for each sentence $u$ and a passage-level hallucination score $\mathcal{H}_{\text{passage}}$ for the entire passage $p$.
@@ 3.2 FactSelfCheck pipeline
As shown in Figure 1, the pipeline of FactSelfCheck consists of three main steps: (1) Knowledge Graph Extraction that extracts sets of entities, relations, and finally, knowledge graph from the initial response $p$ and samples $S$; (2) Fact-level Hallucination Scores, that score facts by measuring factual consistency between facts in $KG_{p}$ and, depending on the variant, $KG_{s}$ in FactSelfCheck-KG or directly $s$ in FactSelfCheck-Text; (3) Sentence-Level and Passage-Level Scores calculation by aggregation of the fact-level scores.
@@ 3.3 Knowledge Graph Extraction
For each instance, we extract a list of entities $\mathcal{E}_{p}$ by passing $p$ to the LLM (Prompt 1).
@@
Next, we provide $p$ and $\mathcal{E}_{p}$ to the LLM to extract relation types between entities, resulting in $\mathcal{R}_{p}$ (Prompt 2).
@@
We then input $\mathcal{E}_{p}$, $\mathcal{R}_{p}$, and each sentence $u\in U$ into the LLM to extract the knowledge graph $KG_{u}$ (Prompt 3). We also provide an initial response $p$ to add contextual information. The output is a set of facts:
@@
After extracting $KG_{u}$ for each sentence $u\in U$, we compile the sets of entities $\mathcal{E}_{S}$ and relations $\mathcal{R}_{S}$ required for extracting knowledge graph $KG_{s}$ from each sample $s$ (Prompt 4) [^2] Although we could theoretically use $\mathcal{E}_{p}$ and $\mathcal{R}_{p}$ for $KG_{s}$ extraction, in practice, LLMs are not sufficiently accurate to extract all entities and relations from the response $p$ when calculating $\mathcal{E}_{p}$ and $\mathcal{R}_{p}$. This results in $KG_{u}$ containing entities and relations not present in $\mathcal{E}_{p}$ and $\mathcal{R}_{p}$, even if the prompt restricts them. Empirical tests showed that extending $\mathcal{E}_{S}$ and $\mathcal{R}_{S}$ by adding entities and relations from all $KG_{u}$ improved the results.:
@@
Extracting $KG_{s}$ by utilizing $\mathcal{E}_{S}$ and $\mathcal{R}_{S}$ is more convenient than extraction without them, as it eliminates the need for entity alignment and ensures that the KG is built using the same schema as $KG_{p}$.
@@ 3.4.1 FactSelfCheck-KG
To allow semantic matching and reasoning over knowledge graphs, rather than only exact fact matching, we introduce the LLM-based fact-level hallucination score. We instruct the LLM to determine whether each fact is supported by the knowledge graphs extracted from the samples (Prompt 5). The LLM is expected to respond with ’yes’ or ’no’. We then average the valid responses from the LLM to get the final score as in Equation (8). Any invalid responses are not included in the averaging.
@@ 3.4.2 FactSelfCheck-Text
In the FactSelfCheck-Text variant, we check if a fact is supported by each textual sample directly without using the knowledge graphs. We prompt the LLM to evaluate whether a fact $f$ is supported by the textual sample $s$ (Prompt 5). As in the previous variant, we average the valid LLM responses using the $\Psi$ function:
@@
3.5 Sentence-Level and Passage-Level Hallucination Scores
While detecting hallucinations at the fact level offers fine-grained insights, there are scenarios where sentence-level or passage-level detection is necessary. To achieve this, we aggregate fact-level scores to compute sentence-level scores $\mathcal{H}_{\text{sentence}}(u)$ and a passage-level score $\mathcal{H}_{\text{passage}}(p)$.
@@ 3.5 Sentence-Level and Passage-Level Hallucination Scores
$\mathcal{H}_{\text{passage}}(p)=\text{Agg}_{u\in U}\mathcal{H}_{\text{sentence}}(u)$ (12)
where $u$ represents a single sentence, and $U$ is the set of sentences of the response $p$. The aggregation function Agg can be based on operations such as $mean$ or $max$. The $mean$ function provides a smoothed factuality score, whereas $max$ is sensitive to the most hallucinated facts or sentences. Additionally, sentences without facts are excluded from the aggregation when calculating the passage-level score.
As discussed in Section 4.1, the dataset choice influences the calculation of the passage-level score in Equation 12. Since the ground truth scores are derived by averaging sentence-level scores, we also employ the $mean$ function for aggregation. It is worth mentioning that a more intuitive approach would be to average the fact-level scores from $KG_{p}$, but such an approach cannot be evaluated with the current dataset. Nevertheless, this technique could be promising for real-world applications and future research.
@@ 4 Experimental Setup
In this section, we describe the experimental setup, including the used data, implementation details, and aspects we investigated.
@@ 4.1 Evaluation Data
We utilized the WikiBio GPT-3 Hallucination Dataset Manakul et al. (2023) 333huggingface.co/datasets/potsawee/wiki_bio_gpt3_hallucination (cc-by-sa-3.0 license) for all experiments. It is designed to evaluate sampling-based hallucination detection methods. It consists of $238$ passage-level annotations and $1908$ sentence-level annotations. As it contains only test data, tuning parameters on this dataset would not be methodologically correct. While the dataset focuses on sentence-level and passage-level evaluation, our approach goes further by providing more fine-grained insights. To ensure a meaningful comparison with SelfCheckGPT, we evaluated these levels using aggregation approaches (see Section 3.5). Following the protocol established by Manakul et al., 2023, we merged the labels major-inaccurate and minor-inaccurate into a single hallucinated class.
Additionally, we conducted a fact-level evaluation to provide a more comprehensive analysis, with the detailed description and results presented in Appendix B.
@@ 4.2 Implementation Details
We employed the Llama-3.1-70B-Instruct model Grattafiori et al. (2024) as the LLM in our method, hosting it locally using vLLM Kwon et al. (2023) on a server with 2xNvidia H100 94GB. The sole exception was the hallucination correction experiment, for which we utilized GPT-4o OpenAI et al. (2024), as detailed in Section 4.5. We set the LLM’s temperature to $0.0$ for all calls, except during hallucination correction (see Section 4.5), where we set it to $0.5$. We implemented the methods and experiments using LangChain Chase (2022) and Hugging Face Datasets Lhoest et al. (2021). All pipeline steps were defined using DVC Kuprieiev et al. (2025) to facilitate reproducibility.
To determine whether a language model’s response was ’yes’ or ’no’ in Equation 9, we parsed the text into individual words and verified the presence of the words ’yes’ or ’no’. If either word was detected, the response was excluded from the averaging process in Equations 8 and 10.
4.3 Evaluation of Sentence-Level and Passage-Level Hallucination Detection
For a fair comparison, we employed the same evaluation protocol as SelfCheckGPT. We reported area under the precision-recall curve (AUC-PR) for the sentence-level task and Pearson and Spearman correlation coefficients for the passage-level task. We evaluated our method against all variants of SelfCheckGPT, including Prompt, NLI, Unigram (Max), QA, and BERTScore. We reproduced the results of the Prompt variant using the same LLM employed in our study, namely Llama-3.1-70B-Instruct, whereas the original method utilized an unspecified release of GPT-3.5-turbo. We obtained the remaining results of SelfCheckGPT from the original paper. We ensured consistency in the evaluation protocol by reviewing their source code 444github.com/potsawee/selfcheckgpt.
4.4 Effect of Sample Size on Detection Performance
We investigated the impact of varying the number of samples on the detection performance of our method. Specifically, we evaluated the performance at both the sentence and passage levels by changing the number of samples from $1$ to $20$. We compared our method with the SelfCheckGPT (Prompt) approach.

### v2

@@ Abstract
Large Language Models (LLMs) frequently generate hallucinated content, posing significant challenges for applications where factuality is crucial. While existing hallucination detection methods typically operate at the sentence level or passage level, we propose FactSelfCheck, a novel black-box sampling-based method that enables fine-grained fact-level detection. Our approach represents text as knowledge graphs consisting of facts in the form of triples. Through analyzing factual consistency across multiple LLM responses, we compute fine-grained hallucination scores without requiring external resources or training data. Our evaluation demonstrates that FactSelfCheck performs competitively with leading sentence-level sampling-based methods while providing more detailed insights. Most notably, our fact-level approach significantly improves hallucination correction, achieving a $35.5\%$ increase in factual content compared to the baseline, while sentence-level SelfCheckGPT yields only a $10.6\%$ improvement. The granular nature of our detection enables more precise identification and correction of hallucinated content. Additionally, we contribute a new dataset for evaluating sampling-based methods – FavaMultiSamples.
@@ 1 Introduction
Figure 1: The pipeline of FactSelfCheck in two variants. For response $p$, entities $\mathcal{E}_{p}$ and relations $\mathcal{R}_{p}$ are extracted, followed by the construction of knowledge graphs $KG_{p}$, for which hallucination scores $\mathcal{H}_{\text{fact}}$ are calculated. Samples’ entities $\mathcal{E}_{S}$ and relations $\mathcal{R}_{S}$ are created by merging $\mathcal{E}_{p}$ and $\mathcal{R}_{p}$ with entities and relations from $KG_{p}$. For each sample $s$, the knowledge graph $KG_{s}$ is extracted. FactSelfCheck-KG assesses the consistency between a fact and all $KG_{s}$. FactSelfCheck-Text assesses the consistency between a fact and all $s$ directly. To obtain sentence-level score $\mathcal{H}_{\text{sentence}}$, fact-level scores are aggregated, as indicated by dashed arrows.
@@
Our granular approach is motivated by two key observations. First, a single sentence can contain multiple facts, with the number of facts varying significantly across sentences, contexts, and domains. This variability makes it challenging to identify hallucinated aspects of generated output precisely when using sentence-level detection. Second, false information can be dispersed throughout a text, as a single fact may appear across multiple sentences. That can mislead the sentence-level detection, as the factuality of a sentence is dependent on the previous sentences Zhang et al. (2024a). These challenges become particularly acute when analyzing long texts generated by LLMs. Fine-grained fact-level detection provides a more precise understanding of text factuality than a sentence analysis. It enables better assessment of content reliability and, as we show later, more effective factuality correction.
@@
We propose FactSelfCheck, a black-box method for fact-level hallucination detection, meaning it does not require access to the model’s internal parameters. This design choice makes our approach universally applicable across any LLM, including closed models, like GPT OpenAI et al. (2024). Following a sampling-based detection paradigm, introduced by Manakul et al., 2023, our method utilizes multiple response generations and analyzes the factual consistency of extracted facts across these samples. This paradigm is based on the phenomenon that factual information remains largely consistent across different generations, while hallucinated content tends to vary or contradict itself between samples Manakul et al. (2023); Wang et al. (2023). This way, we can effectively identify hallucinated facts without relying on external resources (zero-resource) or access to the model’s internal parameters (black-box). Moreover, our method is non-parametric, as it does not require any training, making it easy to implement in any domain without the need for training data. The FactSelfCheck pipeline consists of three main steps: knowledge graph extraction, which extracts sets of facts from the initial response and samples; fact-level hallucination scoring; and calculating sentence-level scores by aggregating fact-level scores.
@@
We evaluated our method using the WikiBio GPT-3 Hallucination Dataset Manakul et al. (2023) and the FavaMultiSamples dataset, a new dataset for evaluating sampling-based methods. We performed sentence-level evaluation by aggregating fact-level scores. Our approach achieves performance comparable to leading sampling-based methods while providing more detailed information about hallucinations. Additionally, we demonstrate that our fact-level approach significantly improves hallucination correction. Compared to a baseline, providing incorrect facts to the correction method leads to a $35.5\%$ increase in factual content, while passing incorrect sentences leads to an $10.6\%$ increase.
@@
    The novel black-box sampling-based method for fact-level hallucination detection – FactSelfCheck. It enables fine-grained hallucination detection in LLM responses without requiring training data or external resources, as it is both non-parametric and zero-resource. We propose two effective approaches for measuring factual consistency across multiple samples: FactSelfCheck-KG using knowledge graph comparisons and FactSelfCheck-Text using direct text comparison.
@@
    The FavaMultiSamples dataset, a novel dataset for evaluating sampling-based methods.
@@
Our code is available on GitHub 111github.com/graphml-lab-pwr/FactSelfCheck license: CC BY-SA 4.0. The FavaMultiSamples dataset is available on Hugging Face 222huggingface.co/datasets/graphml-lab-pwr/FavaMultiSamples license: CC BY-SA 4.0. We publish all the code and data, allowing for the reproduction of the results.
@@ 2 Related work
Black-box approaches operate without access to the model’s internal states and aim to detect hallucinations based solely on the text generated by LLMs. Some of these methods use external resources to collect evidence (Min et al., 2023; Chern et al., 2023). Others leverage LLMs to detect hallucinations like CoVe Dhuliawala et al. (2024), which utilizes the chain-of-thought paradigm for detection. Another category is sampling-based methods, such as SelfCheckGPT (Manakul et al., 2023), which evaluate factuality by generating multiple responses (stochastic samples) and assessing consistency between the original response and these samples. The paradigm of utilizing LLM to check it’s own responses was widely studied and adopted in many works (Kadavath et al., 2022; Lin et al., 2024; Ferraz et al., 2024; Zhang et al., 2024b; Miao et al., 2023). Many of these approaches employ a multi-step decomposition strategy to break down the complex task of hallucination detection into more manageable subtasks, a methodology we also adopt in our approach.
@@ 3 Method
We propose FactSelfCheck, a black-box sampling-based method for fact-level hallucination detection, as illustrated in Figure 1.
@@ 3.1 Notation
Our objective is to compute a fact-level hallucination score $\mathcal{H}_{\text{fact}}$ for each fact $f$ in $KG_{p}$. Subsequently, to facilitate comparisons with other methods, we aggregate these scores to obtain a sentence-level hallucination score $\mathcal{H}_{\text{sentence}}$ for each sentence $u$.
@@ 3.2 FactSelfCheck pipeline
As shown in Figure 1, the pipeline of FactSelfCheck consists of three main steps: (1) Knowledge Graph Extraction that extracts sets of entities, relations, and finally, knowledge graph from the initial response $p$ and samples $S$; (2) Fact-level Hallucination Scoring, that score facts by measuring factual consistency between facts in $KG_{p}$ and, depending on the variant, $KG_{s}$ in FactSelfCheck-KG or directly $s$ in FactSelfCheck-Text; (3) Sentence-Level Score calculation by aggregation of the fact-level scores.
@@ 3.3 Knowledge Graph Extraction
For each instance, we extract a list of entities $\mathcal{E}_{p}$ by passing $p$ to the LLM.
@@
Next, we provide $p$ and $\mathcal{E}_{p}$ to the LLM to extract relation types between entities, resulting in $\mathcal{R}_{p}$.
@@
We then input $\mathcal{E}_{p}$, $\mathcal{R}_{p}$, and each sentence $u\in U$ into the LLM to extract the knowledge graph $KG_{u}$. We also provide an initial response $p$ to add contextual information. The output is a set of facts:
@@
After extracting $KG_{u}$ for each sentence $u\in U$, we compile the sets of entities $\mathcal{E}_{S}$ and relations $\mathcal{R}_{S}$ required for extracting knowledge graph $KG_{s}$ from each sample $s$  [^3] Although we could theoretically use $\mathcal{E}_{p}$ and $\mathcal{R}_{p}$ for $KG_{s}$ extraction, in practice, LLMs are not sufficiently accurate to extract all entities and relations from the response $p$ when calculating $\mathcal{E}_{p}$ and $\mathcal{R}_{p}$. This results in $KG_{u}$ containing entities and relations not present in $\mathcal{E}_{p}$ and $\mathcal{R}_{p}$, even if the prompt restricts them. Empirical tests showed that extending $\mathcal{E}_{S}$ and $\mathcal{R}_{S}$ by adding entities and relations from all $KG_{u}$ improved the results.:
@@
Extracting $KG_{s}$ by utilizing $\mathcal{E}_{S}$ and $\mathcal{R}_{S}$ is more convenient and robust than extraction without them, as it eliminates the need for entity alignment and ensures that the KG is built using the same schema as $KG_{p}$.
@@ 3.4.1 FactSelfCheck-KG
To allow semantic matching and reasoning over knowledge graphs, rather than only exact fact matching, we introduce the LLM-based fact-level hallucination score. We instruct the LLM to determine whether each fact is supported by the knowledge graphs extracted from the samples. The LLM is expected to respond with ’yes’ or ’no’. We then average the valid responses from the LLM to get the final score as in Equation (8). Any invalid responses are not included in the averaging.
@@ 3.4.2 FactSelfCheck-Text
In the FactSelfCheck-Text variant, we check if a fact is supported by each textual sample directly without using the knowledge graphs. We prompt the LLM to evaluate whether a fact $f$ is supported by the textual sample $s$. As in the previous variant, we average the valid LLM responses using the $\Psi$ function:
@@
3.5 Sentence-Level Hallucination Score
While detecting hallucinations at the fact level offers fine-grained insights, there are scenarios where sentence-level detection is necessary. To achieve this, we aggregate fact-level scores to compute sentence-level scores $\mathcal{H}_{\text{sentence}}(u)$.
@@ 3.5 Sentence-Level Hallucination Score
where $u$ represents a single sentence, and $U$ is the set of sentences of the response $p$. The aggregation function Agg can be based on operations such as $mean$ or $max$. The $mean$ function provides a smoothed factuality score, whereas $max$ is sensitive to the most hallucinated facts or sentences.
@@ 4 Experimental Setup
In this section, we describe the experimental setup, including the used data, implementation details, and aspects we investigated. Additionally, we conducted a fact-level evaluation to provide a more comprehensive analysis. The detailed description of it is provided in Appendix C.
@@ 4.1 Evaluation Data
Since the method is designed for detecting hallucinations in long generated passages, finding appropriate datasets was challenging. We evaluated our method two datasets. The first one is the WikiBio GPT-3 Hallucination Dataset Manakul et al. (2023) 444huggingface.co/datasets/potsawee/wiki_bio_gpt3_hallucination (cc-by-sa-3.0 license) (later referred to as WikiBio). To the best of our knowledge, this was the only dataset specifically designed for evaluating sampling-based hallucination detection methods. However, the limited availability of such datasets could hinder the development and evaluation of these approaches. To address this limitation, we developed FavaMultiSamples, a novel dataset specifically designed for evaluating methods that analyze multiple samples. We built it upon the FAVA dataset Mishra et al. (2024) and the detailed description of creation is provided in Appendix A. The detailed statistics of the datasets are provided in Appendix B.
Both datasets contain only test data, making it methodologically incorrect to tune parameters, including classification thresholds, on these datasets. While the datasets focus on sentence-level evaluation, our approach provides more fine-grained insights through fact-level analysis. To ensure a meaningful comparison with SelfCheckGPT, we evaluated these levels using aggregation approaches (see Section 3.5). For WikiBio, we followed the protocol established by Manakul et al., 2023, merging the labels major-inaccurate and minor-inaccurate into a single hallucination class.
4.2 Baseline Models
[…]
We include the two best-performing variants of SelfCheckGPT: Prompt and NLI. We implemented the Prompt variant using the same LLM employed in our method, namely Llama-3.1-70B-Instruct, whereas the original method utilized an unspecified release of GPT-3.5-turbo. For the NLI variant, we used the same model as was used in the original paper555huggingface.co/potsawee/deberta-v3-large-mnli license: Apache license 2.0. Finally, the AttentionScore leverages attention maps (Sriramanan et al., 2024), which is, to the best of our knowledge, the only unsupervised internal state-based method. Its unsupervised nature was crucial since the available datasets do not provide training data. We adapted the AttentionScore for sentence-level detection by implementing two variants: (1) Absolute, which analyzes the complete attention map from the input start to the end of the given sentence, and (2) Relative, which analyzes on the attention map between the start and end of the given sentence.
The AttentionScore and probability-based baselines are white-box methods, requiring access to the model. Since we cannot access all models used to generate the datasets, we passed the sequence of tokens to the proxy LLM to obtain the scores. The proxy LLM was the same as the one used in our method.
@@ 4.3 Implementation Details
We employed the Llama-3.1-70B-Instruct model Grattafiori et al. (2024) as the LLM in all steps of our method and for baselines requiring access to LLM. We hosted it locally using vLLM Kwon et al. (2023) on a server with 2xNvidia H100 94GB. For hallucination correction, we utilized GPT-4o OpenAI et al. (2024), as motivated in Section 4.5. We set the LLM’s temperature to $0.0$ for all calls, except during hallucination correction (see Section 4.5), where we set it to $0.5$. The used prompts are available in the code repository. We implemented the methods and experiments using LangChain Chase (2022), Hugging Face Transformers Wolf et al. (2020) and Hugging Face Datasets Lhoest et al. (2021). All pipeline steps were defined using DVC Kuprieiev et al. (2025) to facilitate reproducibility.
To determine whether a language model’s response was ’yes’ or ’no’ in Equation 9, we parsed the text into individual words and verified the presence of the words ’yes’ or ’no’. If either word was detected, the response was excluded from the averaging process in Equations 8 and 10. Moreover, our pipeline is vulnerable to not detecting facts in short, uninformative sentences; for these cases, we set the score to $0.5$. The analysis of the number of such sentences is provided in Appendix D.3.
4.4 Evaluation of Sentence-Level Hallucination Detection
For a fair comparison, we employed the same evaluation protocol as SelfCheckGPT. We reported area under the precision-recall curve (AUC-PR). We ensured consistency in the evaluation protocol by reviewing their source code 666github.com/potsawee/selfcheckgpt.